OpenStudy (anonymous):

What is the distance between points (-6, 4) and (3, -4)?

OpenStudy (anonymous):

\[\sqrt{(9^{2})+(8^{2})}=\sqrt{81+64}=\sqrt{145}\]

OpenStudy (anonymous):

@Fazeelayaz thank you